# Sufficient digits and density estimation: A Bayesian nonparametric approach using generalized finite Pólya trees

### v1

@@ Appendix A Proofs for Section 3
×$\times\prod_{\begin{subarray}{c}x_{1:n}\in\{0,1\}^{n}:\\ d_{x_{1:n}}=0,\,i_{x_{1:n}}=1\end{subarray}}\frac{\ell_{x_{1:n},0}^{N_{x_{1:n},0}}\ell_{x_{1:n},1}^{N_{x_{1:n},1}}}{\ell_{x_{1:n}}^{N_{x_{1:n},0}+N_{x_{1:n},1}}}\frac{(N_{x_{1:n},0}+N_{x_{1:n},1})^{N_{x_{1:n},0}+N_{x_{1:n},1}-1/2}}{{\sqrt{2\pi}}\,N_{x_{1:n},0}^{N_{x_{1:n},0}-1/2}N_{x_{1:n},1}^{N_{x_{1:n},1}{-1/2}}}$
×$\times\prod_{\begin{subarray}{c}x_{1:n}\in\{0,1\}^{n}:\\ d_{x_{1:n}}=1,\,i_{x_{1:n}}=1\end{subarray}}\frac{\ell_{x_{1:n},0}^{N_{x_{1:n},0}}\ell_{x_{1:n},1}^{N_{x_{1:n},1}}}{\ell_{x_{1:n}}^{N_{x_{1:n},0}+N_{x_{1:n},1}}}\frac{(N_{x_{1:n},0}+N_{x_{1:n},1})^{N_{x_{1:n},0}+N_{x_{1:n},1}-1/2}}{{\sqrt{2\pi}}\,N_{x_{1:n},0}^{N_{x_{1:n},0}-1/2}N_{x_{1:n},1}^{N_{x_{1:n},1}{-1/2}}}.$
In a similar way as (C7) was obtained, we see that each term in the first product above is $O((N_{x_{1:n}}/{(2\pi)})^{1/2})$. Consider any $x_{1:n}$ with $d_{x_{1:n}}=1$ and $i_{x_{1:n}}=1$. Defining
@@
then $\epsilon_{x_{1:n}}\in(0,1)$. By (C3),
@@
×$\qquad\qquad\times\epsilon_{x_{1:n},0}^{1/2}(1-\epsilon_{x_{1:n},0})^{1/2}N_{x_{1:n}}^{1/2}/\sqrt{2\pi}$
which is $O(m^{1/2})$ if $\ell_{x_{1:n},0}/\ell_{x_{1:n}}=\epsilon_{x_{1:n},0}$ and converges exponentially fast to 0 otherwise. To prove (C2), since for $\tilde{n}\in\mathcal{N}$ with $\tilde{n}<n^{*}$,
@@
$\mathrm{P}^{*}(Z_{i}\in I_{x_{1:n}}$ $\,|\,Z_{i}\in I_{x_{1:\tilde{n}}})$
$=\mathrm{P}^{*}(Z_{i}\in I_{x_{1:n}}\,|\,Z_{i}\in I_{x_{1:(n-1)}})\cdots\mathrm{P}^{*}(Z_{i}\in I_{x_{1:(\tilde{n}+1)}}\,|\,Z_{i}\in I_{x_{1:\tilde{n}}})$
@@
×$\hskip 56.9055pt\times\frac{1}{\prod_{j=n^{*}+1}^{n}\prod_{x_{1:j}\in\{0,1\}^{j}}\mathrm{b}(\alpha_{x_{1:j},0}+\tilde{N}_{x_{1:j},0},\alpha_{x_{1:j},1}+\tilde{N}_{x_{1:j},1})}.$
@@
×$\hskip 56.9055pt\times\frac{1}{\prod_{j=\bar{n}+1}^{n}\prod_{x_{1:j}\in A_{j}}\mathrm{b}(\alpha_{x_{1:j},0}+\tilde{N}_{x_{1:j},0},\alpha_{x_{1:j},1})}$
×$\hskip 56.9055pt\times\frac{1}{\prod_{j=\bar{n}+1}^{n}\prod_{x_{1:j}\in\{0,1\}^{j}\setminus A_{j}}\mathrm{b}(\alpha_{x_{1:j},0},\alpha_{x_{1:j},1})}.$
@@
where the term []delimited-[][$\cdots$][ ⋯ ] goes to 00 as $m\rightarrow\infty$, since for some $x_{1:\bar{n}}$, $\tilde{N}_{x_{1:\bar{n}}}\rightarrow\infty$ as $m\rightarrow\infty$. ∎
@@
To prove the statement it is sufficient to show that $f^{*}$ belongs to the Kullback-Leibler support of the prior. This follows from Theorem 6.16 and Example 6.20 in [14].
@@
for any $\bar{n}\in\mathbb{N}_{0}$. Following closely the proof of Theorem 7.1 in [14], we will show that
@@
Again from Theorem 7.1 in [14] we see that for any $\delta>0$,
@@
is negligible by the assumption $\int f^{*}(x)\log f^{*}(x)\,\mathrm{d}x<\infty$ and Lemma B.10 in [14]. Hence, $f^{*}$ belongs to the Kullback-Leibler support of $\Pi_{1}$, whereby the proof is complete in case of the GFPT1 prior.
The proof under the GFPT2 prior follows by observing that conditioned on $R$, the GFPT2 prior coincides with a GFPT1 prior, and then applying Proposition 6.28 in [14]. ∎
For the proof of Theorem 5.9, we need some additional notation. For a (semi)metric space $(S,d)$ and a totally bounded subset $C\subset S$, define the $\varepsilon$-covering number of $C$ as the smallest integer of balls of radius $\varepsilon$ (with respect to $d$) needed to cover $C$. Denote this $\varepsilon$-covering number by $N(\varepsilon,C,d)$. The proof of Theorem 5.9 is based on the following general result, where $\mathcal{P}$ denotes the set of absolutely continuous probability distributions on $[0,1)$ and this set is equipped with the Borel $\sigma$-algebra generated by the Hellinger distance.
Theorem D.2 (Theorem 6.23 in [14]).
@@
We apply Theorem D.2. Since $f^{*}$ belongs to the Kullback-Leibler support of the prior, cf. the proof of Theorem 5.8, we just need to verify items $(i)$ and $(ii)$ in Theorem D.2. To do so, we follow closely the proof of Theorem 7.16 in [14]: For a given PDF $f$ on $[0,1)$, denote by $f_{k}$ its discretization at level $k$ of the NBP $I$, cf. (D9). Consider partitions
@@
such that $A\log m\leq k_{m}\leq B\log m$ where $0<A<B$ are constants. Then the proof of item $(i)$ above follows the same lines as the proof of Theorem 7.16 in [14]. Concerning item $(ii)$, write

### v2

@@ Acknowledgements
$\times\prod_{\begin{subarray}{c}x_{1:n}\in\{0,1\}^{n}:\\ d_{x_{1:n}}=0,\,i_{x_{1:n}}=1\end{subarray}}\frac{\ell_{x_{1:n},0}^{N_{x_{1:n},0}}\ell_{x_{1:n},1}^{N_{x_{1:n},1}}}{\ell_{x_{1:n}}^{N_{x_{1:n},0}+N_{x_{1:n},1}}}\frac{(N_{x_{1:n},0}+N_{x_{1:n},1})^{N_{x_{1:n},0}+N_{x_{1:n},1}-1/2}}{{\sqrt{2\pi}}\,N_{x_{1:n},0}^{N_{x_{1:n},0}-1/2}N_{x_{1:n},1}^{N_{x_{1:n},1}{-1/2}}}$
$\times\prod_{\begin{subarray}{c}x_{1:n}\in\{0,1\}^{n}:\\ d_{x_{1:n}}=1,\,i_{x_{1:n}}=1\end{subarray}}\frac{\ell_{x_{1:n},0}^{N_{x_{1:n},0}}\ell_{x_{1:n},1}^{N_{x_{1:n},1}}}{\ell_{x_{1:n}}^{N_{x_{1:n},0}+N_{x_{1:n},1}}}\frac{(N_{x_{1:n},0}+N_{x_{1:n},1})^{N_{x_{1:n},0}+N_{x_{1:n},1}-1/2}}{{\sqrt{2\pi}}\,N_{x_{1:n},0}^{N_{x_{1:n},0}-1/2}N_{x_{1:n},1}^{N_{x_{1:n},1}{-1/2}}}.$
In a similar way as (C7) in this supplementary material was obtained, we see that each term in the first product above is $O((N_{x_{1:n}}/{(2\pi)})^{1/2})$. Consider any $x_{1:n}$ with $d_{x_{1:n}}=1$ and $i_{x_{1:n}}=1$. Defining
@@
then $\epsilon_{x_{1:n}}\in(0,1)$. By (C3) in this supplementary material,
@@
$\qquad\qquad\times\epsilon_{x_{1:n},0}^{1/2}(1-\epsilon_{x_{1:n},0})^{1/2}N_{x_{1:n}}^{1/2}/\sqrt{2\pi}$
which is $O(m^{1/2})$ if $\ell_{x_{1:n},0}/\ell_{x_{1:n}}=\epsilon_{x_{1:n},0}$ and converges exponentially fast to 0 otherwise. To prove (C2) in this supplementary material, since for $\tilde{n}\in\mathcal{N}$ with $\tilde{n}<n^{*}$,
@@
[…]
@@
$\hskip 56.9055pt\times\frac{1}{\prod_{j=n^{*}+1}^{n}\prod_{x_{1:j}\in\{0,1\}^{j}}\mathrm{b}(\alpha_{x_{1:j},0}+\tilde{N}_{x_{1:j},0},\alpha_{x_{1:j},1}+\tilde{N}_{x_{1:j},1})}.$
@@
$\hskip 56.9055pt\times\frac{1}{\prod_{j=\bar{n}+1}^{n}\prod_{x_{1:j}\in A_{j}}\mathrm{b}(\alpha_{x_{1:j},0}+\tilde{N}_{x_{1:j},0},\alpha_{x_{1:j},1})}$
$\hskip 56.9055pt\times\frac{1}{\prod_{j=\bar{n}+1}^{n}\prod_{x_{1:j}\in\{0,1\}^{j}\setminus A_{j}}\mathrm{b}(\alpha_{x_{1:j},0},\alpha_{x_{1:j},1})}.$
@@
where the term $\cdots$] goes to 0 as $m\rightarrow\infty$, since for some $x_{1:\bar{n}}$, $\tilde{N}_{x_{1:\bar{n}}}\rightarrow\infty$ as $m\rightarrow\infty$. ∎
@@
To prove the statement it is sufficient to show that $f^{*}$ belongs to the Kullback-Leibler support of the prior. This follows from Theorem 6.16 and Example 6.20 in Ghoshal and van der Vaart (2017).
@@
for any $\bar{n}\in\mathbb{N}_{0}$. Following closely the proof of Theorem 7.1 in Ghoshal and van der Vaart (2017), we will show that
@@
Again from Theorem 7.1 in Ghoshal and van der Vaart (2017) we see that for any $\delta>0$,
@@
is negligible by the assumption $\int f^{*}(x)\log f^{*}(x)\,\mathrm{d}x<\infty$ and Lemma B.10 in Ghoshal and van der Vaart (2017). Hence, $f^{*}$ belongs to the Kullback-Leibler support of $\Pi_{1}$, whereby the proof is complete in case of the GFPT1 prior.
The proof under the GFPT2 prior follows by observing that conditioned on $R$, the GFPT2 prior coincides with a GFPT1 prior, and then applying Proposition 6.28 in Ghoshal and van der Vaart (2017). ∎
For the proof of Theorem 5.9, we need some additional notation. For a (semi)metric space $(S,d)$ and a totally bounded subset $C\subset S$, define the $\varepsilon$-covering number of $C$ as the smallest integer of balls of radius $\varepsilon$ (with respect to $d$) needed to cover $C$. Denote this $\varepsilon$-covering number by $N(\varepsilon,C,d)$. The proof of Theorem 5.9 is based on the following general result (Theorem 6.23 in Ghoshal and van der Vaart, 2017) where $\mathcal{P}$ denotes the set of absolutely continuous probability distributions on $[0,1)$ and this set is equipped with the Borel $\sigma$-algebra generated by the Hellinger distance.
Theorem D.2.
@@
We apply Theorem D.2. Since $f^{*}$ belongs to the Kullback-Leibler support of the prior, cf. the proof of Theorem 5.8, we just need to verify items $(i)$ and $(ii)$ in Theorem D.2. To do so, we follow closely the proof of Theorem 7.16 in Ghoshal and van der Vaart (2017): For a given PDF $f$ on $[0,1)$, denote by $f_{k}$ its discretization at level $k$ of the NBP $I$, cf. (D9). Consider partitions
@@
such that $A\log m\leq k_{m}\leq B\log m$ where $0<A<B$ are constants. Then the proof of item $(i)$ above follows the same lines as the proof of Theorem 7.16 in Ghoshal and van der Vaart (2017). Concerning item $(ii)$, write